### v1

# Learning from Interaction: A Survey

Abstract
LLM-based agents can autonomously accomplish complex tasks across various domains. However, to further cultivate capabilities such as adaptive behavior and long-term decision-making, training on static datasets built from human-level knowledge is insufficient. These datasets are costly to construct and lack both dynamism and realism. A growing consensus is that agents should instead interact directly with environments and learn from experience through reinforcement learning. We formalize this iterative process as the Generation-Execution-Feedback (GEF) loop, where environments generate tasks to challenge agents, return observations in response to agents’ actions during task execution, and provide evaluative feedback on rollouts for subsequent learning. Under this paradigm, environments function as indispensable producers of experiential data, highlighting the need to scale them toward greater complexity, realism, and interactivity. In this survey, we systematically review representative methods for environment scaling from a pioneering environment-centric perspective and organize them along the stages of the GEF loop, namely task generation, task execution, and feedback. We further analyze benchmarks, implementation strategies, and applications, consolidating fragmented advances and outlining future research directions for agent intelligence. [^1] We provide a GitHub repository with real-time updates on this topic: https://github.com/lukahhcm/Awesome_Scaling_Environments.
Scaling Environments for LLM Agents in the Era of
Learning from Interaction: A Survey
Yuchen Huang Sijia Li Minghao Liu Wei Liu Shijue Huang Zhiyuan Fan Hou Pong Chan Yi R. (May) Fung Hong Kong University of Science and Technology King’s College London  DAMO Academy, Alibaba Group {yhuanggn, yrfung}@cse.ust.hk
Figure 1: (A) Experience arises from the Generation-Execution-Feedback (GEF) loop, where environments generate tasks, agents execute them, and environments evaluate and filter useful experience for RL training. (B) Overview of environment scaling: a GEF-aligned taxonomy of environment-scaling methods, alongside implementation, applications, and the unique challenge of Generator-Verifier asymmetry.
1 Introduction
The rapid progress of large language models (LLMs) has catalyzed a transformative shift in artificial intelligence, precipitating a surge of research on LLM-based agents (luo2025largelanguagemodelagent; xi2025rise). Such agents inherit strong reasoning and task-decomposition capabilities from their base models and, when augmented with modules for tool use and memory, can execute actions, interact with real or simulated environments, accumulate experience over time, and progressively improve their own behavior. This design has achieved remarkable progress across diverse domains, including automated coding (qwen3technicalreport; anthropic2025claude4), interactive web navigation (openai2025introducing; he2025advancing), tool use (zhang2025agentorchestra; anthropic2024mcp), and deep research (tongyidr; openai2025deepresearch; google2024geminideepresearch).
However, as agent capabilities continue to evolve, it is infeasible to attain intelligence beyond the human-level merely by supervised fine-tuning (SFT) pretrained models on static datasets (huang2025r; su2025crossingrewardbridgeexpanding; zhao2025absolutezeroreinforcedselfplay). Such datasets are typically manually annotated or curated under human oversight, which makes them costly and labor-intensive to produce at scale, intrinsically bounded by human-level knowledge, and lacking realism and adaptability. By contrast, reinforcement learning provides a more aligned training paradigm (tao2024surveyselfevolutionlargelanguage; zhang2025landscapeagenticreinforcementlearning), where agents can explore in the environment, accumulate experiences, and finally acquire new knowledge or skills. We formalize this interactive process as the Generation-Execution-Feedback (GEF) loop, illustrated in Figure 1 (A). In each iteration, the environment first generates diverse tasks, then the agent executes them within the environment, producing action-observation trajectories. The environment subsequently evaluates these rollouts and retains useful experience for subsequent training. Repeated iterations progressively refine the policy and expand the agent’s capabilities. Notably, unlike prior work (gao2025survey), we adopt a broad view of the environment: everything external to the current agent, including the state space, the executable action space, the design of feedback for interaction and evaluation, and the activities of users and other agents, is considered part of it. In this setting, the environment is no longer a mere container for agents’ activities; it has become an active producer of experiential data, underscoring the growing need for scaling environments to create a more complex, realistic, and richly interactive world (camel2025scaling).
Recent research has embraced this trend of scaling the environment from different perspectives. For instance, systems like AgentGen (hu2025agentgen), AgentGym (xi2024agentgym), and GEM (liu2025gemgymagenticllms) devise heterogeneous environments to increase the diversity of the generated tasks. R-Zero (huang2025r) proposes a challenger-solver framework that autonomously generates increasingly difficult tasks. RandomWorld (sullivan2025procedural) scales up the interactivity by procedural generation of diverse tools for agents to access. ARE (andrews2025arescalingagentenvironments) develops an event-driven environment that supports asynchronous interactions between the environment and agents, scaling up the environmental dynamics that conform to realistic settings. However, a systematic analysis that connects these research directions remains absent.
Scaling Environments for Agents
Stage 1: Task Generation (§3) Complexity (§3.1) e.g., ToolLLM (qin2023toolllm), BFCL V3 (patil2025bfcl), $\tau$-bench (yao2024tau), $c^{3}$-bench (yu2025c), TaskCraft (shi2025taskcraft), WebDancer (wu2025webdancerautonomousinformationseeking), WebWalker (wu2025webwalkerbenchmarkingllmsweb), WebSailor (li2025websailornavigatingsuperhumanreasoning), WebShaper (tao2025webshaperagenticallydatasynthesizing), etc. Dynamic  (§3.2) e.g., Eurekaverse (liang2024environment), EvoCurr (cheng2025evocurr), WebRL (qiwebrl), AgentGym (xi2024agentgym), AgentGen (hu2025agentgen), R-Zero (huang2025r), EnvGen (zala2024envgen), ARE (andrews2025arescalingagentenvironments), etc. Diversity  (§3.3) e.g., AgentGym (xi2024agentgym), AgentSense (leng2025agentsense), AgentBank (song2024agentbank), etc. Stage 2: Task Execution (§4) Interactivity (§4.1) e.g., RandomWorld (sullivan2025procedural), AppWorld (trivedi2024appworldcontrollableworldapps), WebShop (yao2023webshopscalablerealworldweb), MCP-Universe (luo2025mcp), MCP-Bench (wang2025mcp), MCPToolBench++ (fan2025mcptoolbench++), AgentScaler (fang2025generalagentic), BrowseMaster (pang2025browsemaster), etc. Realism (§4.2) e.g., RestGPT (song2023restgpt), WebWalker (wu2025webwalkerbenchmarkingllmsweb), AutoMCP (mastouri2025makingrestapisagentready), Tongyi DeepResearch (tongyidr), AgentScaler (fang2025generalagentic), FTRL (ye2025feedbackdriven), $\tau$-bench (yao2024tau), $\tau^{2}$-bench (barres2025tau2bench), APIGen-MT (prabhakar2025apigenmt), ToolHop (ye2025toolhop), Genie 3 (parker2025genie3), Oasis (yang2025oasis), AgentSociety (zhang2025parallelized), ARE (andrews2025arescalingagentenvironments), etc. Stage 3: Feedback  (§5) Density  (§5.1) e.g., VerlTool (jiang2025verltool), Agent-RLVR (da2025agent), Toolrl (qian2025toolrl), SRM (ma2025steplevelrewardmodelsrewarding), ThinkPRM (khalifa2025processrewardmodelsthink), Web-Shepherd (chae2025webshepherd), PR-Clip-Delta (gao2024designing), etc. Granularity  (§5.2) e.g., AdaCtrl (huang2025adactrl), Toolrl (qian2025toolrl), R1-Searcher (song2025r1), RaR (gunjal2025rubricsrewardsreinforcementlearning), Rubicon (huang2025reinforcement), RuscaRL (zhou2025breaking), etc. Automation (§5.3) e.g., JudgeLRM (chen2025judgelrmlargereasoningmodels), GenRM-CoT (zhang2025generativeverifiersrewardmodeling), REWARDAGENT (peng2025agenticrewardmodelingintegrating), ARMAP (chen2025scalingautonomousagentsautomatic), RaR (gunjal2025rubricsrewardsreinforcementlearning), Rubicon (huang2025reinforcement), etc. Objectivity (§5.4) e.g., RM-7B (su2025crossingrewardbridgeexpanding), General-Reasoner (ma2025general), Rlpr (yu2025rlpr), Nemotron-Crossthink (akter2025nemotron), Nover (liu2025nover), Writing-Zero (jia2025writingzerobridgegapnonverifiable), Omni-Thinker (li2025omni), ARE (andrews2025arescalingagentenvironments), etc. Robustness (§5.5) e.g., ScoreDiff (lin2024navigating), MONA (farquhar2025monamyopicoptimizationnonmyopic), PAR (fu2025reward), RRM (liu2024rrm), InfoRM (miao2024inform), RRM (liu2024rrm), RM-7B (su2025crossingrewardbridgeexpanding), Rubicon (huang2025reinforcement), Trinity-RFT (pan2025trinityrftgeneralpurposeunifiedframework), Tongyi DeepResearch (tongyidr), etc.
Figure 2: GEF-aligned taxonomy of environment scaling with dimensions for Task Generation, Task Execution, and Feedback. Representative works are illustrated as leaves on the branches.
Therefore, we comprehensively investigate current environment scaling methods and propose a unified taxonomy aligned with the stages of the GEF loop, adopting a pioneering environment-centric perspective. In the task generation stage, we categorize scaling methods into complexity scaling, dynamic scaling, and diversity scaling, which together characterize an environment’s ability to generate challenging, adaptive, and diverse tasks continuously. In the task execution stage, we highlight interactivity and realism, since these properties determine the richness and fidelity of the interaction data from which agents learn. In the feedback stage, we categorize the scaling of evaluative signals along density, granularity, automation, objectivity, and robustness. Beyond this taxonomy, we also analyze current evaluation benchmarks, implementation frameworks, applications, and future research directions. Figure 1 (B) shows a high-level overview of environment scaling, and representative works are listed in Figure 2.
The survey is organized as follows. We first introduce the background and conceptual framework in §2 and §A. We then categorize representative environment scaling methods along the three-stage taxonomy: task generation (§3), task execution (§4), and feedback (§5). Next, we discuss evaluation benchmarks in §B, implementation frameworks in §6, and applications in §C. Finally, we outline future research directions (§7).
2 Background
Scaling Laws for LLM Agents
Just as large language models exhibit predictable performance scaling with increases in the number of parameters, the volume of training data, and the compute budget, agent systems likewise display scaling regularities along three axes: (i) expanding the agent population and identifying properties that emerge as interactions increase; (ii) increasing environmental complexity and assessing how realistic, dynamic settings shape learning and adaptation; and (iii) extending the horizons of evolution and memory to study how agents generalize and improve through accumulated experience (camel2025scaling). While most existing surveys on LLM agents adopt an agent-centric view (luo2025largelanguagemodelagent; xi2025rise; yehudai2025survey; gao2025survey), covering topics from multi-agent interaction (qian2024scaling; tran2025multi) to self-evolution (gao2025survey; tao2024surveyselfevolutionlargelanguage), environment scaling remains underexplored and has not been systematically organized. In this work, we take an environment-centric perspective on scaling environments and examine how dynamic, richly interactive, high-fidelity worlds can accelerate agent development and evolution.
Generator-Verifier Asymmetry Challenge
A fundamental characteristic in many real-world tasks is the inherent Generator-Verifier Asymmetry (wei2025asymmetry), namely the mismatch between the intelligence required for generator, which generates (§3) or executes (§4) tasks, and that required for verifier, which provides feedback (§5). These two kinds of intelligence naturally form two axes critical to next-generation Agentic AI, as illustrated in Figure 1 (B). From this perspective, scaling up environments essentially corresponds to scaling intelligence along the $x$-axis and the $y$-axis. Current progress in RL largely exploits the regime on the easy-to-verify side of this asymmetry. These Easy-to-Verify, Hard-to-Generate Domains include fields such as mathematics and programming (wei2025browsecomp; jimenez2023swe; phan2025humanity). For these domains, generating and solving a continual stream of high-quality, non-trivial tasks is challenging. In contrast, verification is objective and computationally inexpensive (e.g., via unit tests or exact match on mathematical results). This enables weak-to-strong supervision, where a simple verifier can provide accurate feedback to train a much stronger agent for solving hard tasks. On the contrary, the Hard-to-Verify, Easy-to-Generate Domains include areas such as creative writing, policy-making, or healthcare (lin2024wildbench; arora2025healthbench). For these easy-to-propose, open-ended tasks, verification is subjective, requires substantial expert judgment, or unfolds over long horizons, making high-quality feedback scarce and expensive. This bottleneck, corresponding to the upper-left region of coordinate system, poses more difficulty in modeling the environment, and rendering environment scaling more challenging yet offering greater potential for advancing agent capabilities. Notably, the asymmetry also presents an opportunity: if the generator’s stronger intelligence can be systematically leveraged to strengthen the verifier, so that it can supervise an even stronger generator, then such asymmetric property can be exploited to drive agents’ self-evolution (huang2025r; hong2025cooper; lu2025urpo; chen2025self; wang2025socraticzerobootstrappingreasoning).
Figure 3: Illustration of environment scaling in the task generation and task execution stages, using the example of conference scheduling. Given a user intent, the environment produces a set of tasks for the agent to complete. Scaling in the task generation stage covers complexity scaling, dynamic scaling, and diversity scaling, while in the task execution stage scaling encompasses interactivity scaling and realism scaling.
3 Stage 1: Task Generation
In the task generation stage, the environment is required to propose challenging tasks that push the agent toward its capability boundary. Scaling at this stage targets three aspects of the task design: increasing difficulty (complexity scaling § 3.1), introducing dynamics (dynamic scaling § 3.2), and expanding diversity (diversity scaling § 3.3). An illustrative example is shown in Figure 3. For clarity, in complexity scaling, we only consider the intrinsic difficulty of a task (i.e., static complexity). We group the temporal evolution of task difficulty (dynamic complexity) together with changes in the environment itself under the dynamic scaling subsection.
3.1 Complexity Scaling
Static complexity increases a task’s inherent structural intricacy, moving beyond single-step commands to challenges defined by dependencies, logical flows, and hierarchical relationships. A typical example is LLM tool use, from early single-step tasks to multi-turn, multi-step scenarios, where complexity scales up as the number of turns and steps increases (qin2023toolllm; patil2025bfcl; yao2024tau; yu2025c). More sophisticated tasks exhibit hierarchical or compositional structure, decomposing high-level objectives into nested sub-goals and thereby testing compositional generalization, namely an agent’s ability to solve novel problems by recombining known skills (shao2023compositional). TaskCraft (shi2025taskcraft) operationalizes this by expanding tasks both in depth (longer sequences of tool executions) and in width (multiple sub-goals per objective), enhancing hierarchical reasoning. At the highest level, conditional and graph-based tasks involve non-linear structures with branching logic, where planning must adapt dynamically to intermediate outcomes. Recent efforts in information-seeking agents extend linear sequences to complex graph-based information chains (wu2025webdancerautonomousinformationseeking; tao2025webshaperagenticallydatasynthesizing; li2025websailornavigatingsuperhumanreasoning; wu2025webwalkerbenchmarkingllmsweb), and multi-agent settings further amplify this complexity, as optimal plans become contingent on the actions of other agents, producing intrinsically interdependent, graph-structured challenges.
3.2 Dynamic Scaling
Task Difficulty Dynamics
Scaling the task’s difficulty dynamically helps agents generalize, which makes the targets non-stationary and changes the actions and states, and we can tune this either on a predetermined schedule or based on how the agent performs. A common strategy is performance-driven scheduling, where the task difficulty is regulated by the success rate (SR), as in Eurekaverse (liang2024environment). Other approaches target newly acquired or weaker skills, as in EvoCurr (cheng2025evocurr) and EnvGen (zala2024envgen). AgentGen’s BI-EVAL mechanism (hu2025agentgen) introduces a bidirectional variation, which adjusts complexity upward or downward to match agent capability, in contrast to earlier methods that mostly increased difficulty (xu2025wizardlmempoweringlargepretrained; luo2025wizardmathempoweringmathematicalreasoning). Beyond these, WebRL (qiwebrl) implements self-adjusting curricula across complex web settings, and AgentGym (xi2024agentgym) generalizes performance-adaptive scheduling to diverse benchmarks. R-Zero (huang2025r) formalizes a challenger-solver paradigm in which a challenger proposes near-boundary tasks based on the solver’s uncertainty, and the solver improves by training on filtered task sets, yielding iterative and targeted curricula.
Environmental Dynamics
Environmental dynamics provide more realistic scenarios for agents. For example, the Meta Agents Research Environments (ARE) platform (andrews2025arescalingagentenvironments) pushes this paradigm further by introducing a more realistic and dynamic environment. In most setups, if the agent is not interacting, the environment typically remains frozen. By contrast, ARE allows the environment to change dynamically through random or scheduled events at all times, allowing it to evolve asynchronously and independently of the agent. As shown in ARE’s Mobile and Gaia2 benchmarks (andrews2025arescalingagentenvironments), agents in such dynamic settings need to balance cognitive depth and temporal responsiveness to manage interaction latency and uncertainty as the environment’s state continually changes. This design shifts the focus from static interaction to more continuous and proactive engagement.
3.3 Diversity Scaling
Scaling the diversity of data is key to building more robust and generalizable agents. Managing diversity at the task level (e.g., task difficulty, task objectives) helps agents acquire broader skills rather than overfitting to specific patterns (hu2025agentgen; huang2025r). At the environment level, exposing agents to a wide range of scenarios (e.g., different domains or tool suites) can further enhance their adaptability to novel situations. For example, representative works like AgentGen (hu2025agentgen) and AgentGym (xi2024agentgym) synthesize a wide range of heterogeneous settings that broaden the training signals agents receive. Beyond these, AgentSense (leng2025agentsense) generates diverse virtual sensor data by simulating different human personas and routines, and AgentBank (song2024agentbank) shows that training on tens of thousands of heterogeneous interaction trajectories will substantially improves generalization. Collectively, these approaches demonstrate that diversity across tasks and environments is foundational for training capable, adaptable agents.
4 Stage 2: Task Execution
In the task execution stage, after the agent takes an action, it receives an observation from the environment. Consequently, whether the agent can interact with the environment in real time (interactivity § 4.1) and whether the returned observations are consistent with real-world scenarios (realism § 4.2) are both critical to the quality of the resulting experience. Accordingly, we organize environment scaling in this stage into two directions: interactivity scaling and realism scaling, as shown next to the action space in Figure 3.
4.1 Interactivity Scaling
Despite the advent of standard protocols such as the Model Context Protocol (MCP) (anthropic2024mcp; luo2025mcp; wang2025mcp; fan2025mcptoolbench++) integrates heterogeneous data sources and tools into a unified form of context and thus greatly improves the efficiency and controllability of tool use, many datasets (fan2025mcptoolbench++; liu2024apigenautomatedpipelinegenerating; qin2023toolllm) still consist of predefined, carefully curated sequences of tool calls, even some of them include real API or MCP calls. Under this non-interactive settings, each task has a single predefined solution path. Agents are blind to intermediate tool outputs and cannot adapt subsequent tool selection based on the returned results. Consequently, agents trained on this static supervision exhibit poor generalization to novel tasks and limited diversity in solution paths (sullivan2025procedural). Recent methods (pang2025browsemaster; wang2024executable; yao2023webshopscalablerealworldweb; trivedi2024appworldcontrollableworldapps; sullivan2025procedural) start to allow agents to interactively invoke real world APIs or leverage tools via function calling or code generation, where they can adjust subsequent tool selections based on current output. Among these approaches, BrowseMaster (pang2025browsemaster) further supports parallel tool calling, expanding the typical one-tool-call-per-invocation pattern to an average of 12.11 calls per invocation, which further increases interactivity. Another promising direction uses an offline real database as the interaction environment (tongyidr; fang2025generalagentic; ye2025feedbackdriven; yao2024tau; barres2025tau2bench; prabhakar2025apigenmt; ye2025toolhop), where the agent interactively calls functions to read and write the state database. This approach strikes a practical balance between interactivity and realism and helps agents accumulate meaningful experience efficiently.
4.2 Realism Scaling
To ensure that large language model (LLM) agents can generalize effectively to complex, real-world scenarios, the training data derived from these environments should maintain real-world consistency. In tool-use environments, earlier works (qin2023toolllm; lu2025toolsandbox; sun2025zerosearch) utilize LLMs to generate the results of tool calls as simulations, avoiding monetary cost and unexpected implementation errors. However, to better ensure real-world consistency, more recent approaches start to use real APIs (song2023restgpt; wu2025webwalkerbenchmarkingllmsweb; mastouri2025makingrestapisagentready) or execute tasks in simulated environments backed by offline, real-world databases (tongyidr; fang2025generalagentic; ye2025feedbackdriven; yao2024tau; barres2025tau2bench; prabhakar2025apigenmt; ye2025toolhop). Specifically, tongyidr builds a custom tool suite and a simulated environment based on an offline Wikipedia database, where agents execute actual tool calls to directly read and write to the database, achieving lower cost and higher efficiency. Beyond tool-use scenarios, where interactions are mainly text-based, the emerging paradigm of thinking with images (su2025thinking) advocates using visual information as a dynamic, manipulable workspace for intermediate reasoning. Genie 3 (parker2025genie3) further extends realism scaling to 3D scenarios by implementing a physically grounded, real-time interactive 3D world. It preserves realistic physical properties while improving long-horizon consistency, offering a practical framework and motivating the development of more realistic environments for agents to explore.
Another path to improving realism of environments is the simulation of multi-agent settings. In such contexts, agents may coordinate or compete with each other (tran2025multi; qian2024scaling; li2024smoa; zhang2024chain; kim2024llmsproducefaithfulexplanations), therefore each agent’s behavior naturally becomes part of the environment for the others. As the number of agents scales, these interactions can produce emergent social and economic phenomena such as information diffusion, opinion polarization, and herding effects (yang2025oasis; zhang2025parallelized). To reproduce such societal dynamics and improve data fidelity, multi-agent frameworks like Oasis (yang2025oasis) leverage real world social media data stored in a relational database to simulate interactive environments, enabling more realistic modeling of social processes. Meanwhile, zhang2025parallelized move beyond earlier frameworks (li2023camel; gao2024agentscope) that enforce execution order via message passing SOPs by adopting the MQTT communication protocol to support asynchronous decision making among autonomous agents, more closely simulating real-world workflows. Similarly, ARE  (andrews2025arescalingagentenvironments) decouples agent and environment clocks so that world state evolves asynchronously, and it simulates other agents’ activities by treating them as independent events in the environment’s event stream. Such environments enable the simulation of real-world societal processes and support the collection of more realistic experiential data.
5 Stage 3: Feedback
Figure 4: Illustration of environment scaling in the feedback stage using a conference-scheduling example. The agent first executes tasks in the environment and produces action-observation trajectories. The environment then evaluates these trajectories and returns feedback, yielding the experience used to train the agent. Scaling in the feedback stage covers density, granularity, automation, objectivity, and robustness.
In the feedback stage, the environment assesses the trajectories collected during task execution and generates feedback signals for subsequent RL training. Scaling at this stage focuses on how feedback is provided, including its frequency and richness (density § 5.1 and granularity § 5.2), its level of automation (automation § 5.3), as well as how objectively and reliably it is delivered (objectivity § 5.4 and robustness § 5.5). Accordingly, we categorize representative scaling approaches along five dimensions: density, granularity, automation, objectivity, and robustness, as shown in Figure 4.
5.1 Density Scaling
The density of feedback refers to how frequently the environment provides evaluative feedback signals. Early environments typically provided trajectory-level rewards based on the final outcome (success or failure) (jiang2025verltool; da2025agent; qian2025toolrl; wang2025otc). Despite such sparse signals leading to more stable training, especially in tasks like mathematics or code generation (deepseekai2025deepseekr1incentivizingreasoningcapability), they are insufficient for more complex multi-step tasks. In these tasks, it is challenging to verify whether an intermediate step contributes to or hinders task completion. To address this issue and scale feedback density, more recent studies (khalifa2025processrewardmodelsthink; chae2025webshepherd; gao2024designing; ma2025steplevelrewardmodelsrewarding; park2025know) have introduced step-level process-based rewards, often combined with traditional trajectory-level outcome-based rewards, thus providing denser supervision and more detailed guidance for agent improvement. However, the mechanism behind such reward designs remains underexplored, though some interesting phenomena have been observed in existing experiments. For instance, ma2025steplevelrewardmodelsrewarding shows that step-level reward models can perform well in logical-coherent tasks such as mathematical reasoning, but they are not suitable for some natural language tasks (e.g., creative writing, policy making). gao2024designing shows that naively combining step-level rewards with trajectory-level rewards can lead to reward hacking. That is, trivial actions may get high rewards, resulting in repetitive reasoning behavior that ultimately harms the agent’s training process. They further use reward differences between adjacent steps as rewards and clip them if they exceed a predefined threshold to mitigate this problem. Apart from these, park2025know finds that even state-of-the-art reward models can be poorly calibrated, assigning overly optimistic scores to some intermediate steps. These studies collectively showcase both the potential and bottlenecks of current dense reward designs. In the future, designing reward models that are more accurately calibrated, stable in training, and interpretable will be the focus of the reward density scaling.
5.2 Granularity Scaling
Granularity scaling refers to increasing the level of detail in feedback as well as enriching the forms of feedback provided. In earlier stages, evaluative feedback typically consisted of binary signals (NIPS2017_d5e2c0ad; ibarz2018rewardlearninghumanpreferences) or a single scalar score (stiennon2022learningsummarizehumanfeedback; ouyang2022traininglanguagemodelsfollow). At finer granularities, the evaluative feedback on agents’ performances is broken down into structured components, such as a set of scores across multiple criteria, offering more informative guidance (huang2025adactrl). Building on this idea, some works use manually designed, multi-faceted rewards to improve correctness and format (qian2025toolrl; song2025r1). The Rubrics as Rewards (RaR) framework (gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement; zhou2025breaking; zhang2025chasingtail; viswanathan2025checklist) further decomposes task requirements into tangible, human-interpretable criteria. By designing rewards as checklist-style, instance-specific rubrics, it provides a middle ground between binary correctness signals and broad preference rankings.
5.3 Automation Scaling
Automation scaling refers to the process where the feedback mechanism shifts from human supervision to automated evaluations generated by artificial intelligence. Reinforcement Learning from Human Feedback (RLHF) encompasses a wide range of frameworks, with the goal of aligning large language models (LLMs) with human preferences (Sheng_2025; hu2025openrlhfeasytousescalablehighperformance; christiano2017deep). Although this method is effective, it heavily relies on human annotators, which results in a slow, costly and difficult-to-scale feedback process. In order to achieve automated feedback and reduce the reliance on labor-intensive preference annotations, an increasing number of studies have begun to use LLMs with evaluation capabilities as automated judges to replace human evaluators, thereby forming a new paradigm called "Reinforcement Learning from AI Feedback" (RLAIF) (zhang2025generativeverifiersrewardmodeling; su2025crossingrewardbridgeexpanding; chen2025judgelrmlargereasoningmodels; lee2024rlaifvsrlhfscaling; bai2022constitutionalaiharmlessnessai). The pioneering work REWARDAGENT (peng2025agenticrewardmodelingintegrating) introduced a verification agent to simultaneously assess the factual correctness of model responses and their compliance with instructions, and combine these assessment results with basic human feedback to effectively guide model training. In contrast, ARMAP (chen2025scalingautonomousagentsautomatic) innovatively bypassed the need for a more powerful LLM to act as a judge, instead using the positive and negative trajectories generated by the LLM to train the classification model, thereby constructing an automatic and efficient reward mechanism. Rubric-based methods (gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement), on the other hand, decompose open-ended tasks into multi-dimensional, interpretable criteria, improving the quality of rewards generated by LLMs. By integrating this automated assessment into the training environment, agents can gradually acquire higher-level capabilities without the need for continuous human supervision. However, this paradigm also magnifies risks such as the propagation of LLM’s own biases in the process of adjudication, as well as the phenomenon of rewarding hacking, which refers to the situation where intelligent agents exploit the loopholes in the reward model to strive for high scores, thereby deviating from the intended goals and even causing the model to crash. This reveals the crucial trade-off between scalability and security, and emphasizes the need for research on robust risk mitigation strategies.
5.4 Objectivity Scaling
The RLVR paradigm has achieved great success in objective and easy-to-verify domains such as mathematical reasoning and code generation (shao2024deepseekmathpushinglimitsmathematical; deepseekai2025deepseekr1incentivizingreasoningcapability). However, in many real-world scenarios such as creative writing, medical consultation, and policy making, the objectivity is subjective, open, and hard to verify. In these real environments, the feedback collected often contains many biases and noise. An environment for training agents should be able to extract accurate and high-quality feedback from such kinds of objectivity, and gradually scale from simple and verifiable tasks to difficult and hard-to-verify ones. Some recent studies have made early attempts in this direction (andrews2025arescalingagentenvironments; ma2025general; akter2025nemotron; su2025crossingrewardbridgeexpanding; yu2025rlpr; liu2025nover; gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement). For example, Writing-Zero (jia2025writingzerobridgegapnonverifiable) uses a pairwise generative model to extract reliable and verifiable signals from subjective evaluations. Omni-Thinker (li2025omni) combines rule-based verifiable rewards with generative preference signals to form a unified multi-task RL training loop. ARE’s verifier (andrews2025arescalingagentenvironments) applies RLVR by using both hard checks (exact parameter matches) and soft checks (LLM-based semantic judgments) under strict causal order. Although these works show the feasibility of structured and verifiable rewards in open-ended environments, scaling environments to handle harder forms of objectivity still faces many challenges.
5.5 Robustness Scaling
Feedback robustness scaling requires the environment to provide more stable and reliable reward signals. Regarding the robustness of the reward itself, signals may be noisy, or face the challenge of reward hacking, which refers to agents learning undesirable and tricky behavioral patterns that obtain high reward without achieving the intended goals (miao2024inform; liu2024rrm; farquhar2025monamyopicoptimizationnonmyopic; fu2025reward; tarek2025reward). For the former issue, some approaches generate soft probabilistic rewards with generative verifiers to mitigate noise in rewards (lin2024navigating; su2025crossingrewardbridgeexpanding). And as for the latter, frameworks such as MONA (farquhar2025monamyopicoptimizationnonmyopic) try to mitigate reward hacking problem by evaluating the future utility of actions through an overseer, thus constraining unstable behaviors while preserving explainability. huang2025reinforcement develops a reward hacking defense rubric that penalizes sycophantic praise towards user prompts and overly flattering self-assessments in responses, encouraging the model to produce more substantive content.
Apart from reward robustness, interaction robustness at environment-level is also crucial, since it is common for environmental instabilities (e.g., delays, crashes, corrupted tool outputs) to undermine feedback reliability and degrade the training process. To handle such failures, Trinity-RFT (pan2025trinityrftgeneralpurposeunifiedframework) proposes asynchronous inference and retry mechanisms, while tongyidr employs caching, retrying failed calls, and switching to similar providers to prevent corrupted trajectories. Looking ahead, future reward design should prioritize both efficacy and the prevention of hacking through tricky patterns, and greater attention should be paid to the design of more robust system.
6 Implementation Frameworks
There exist diverse implementations of simulated environments that vary in modality and complexity.
Visual Environments
For 2D scenarios, tasks in grid-based environments can be symbolic or pixel-level, often associated with game-playing (chevalier2023minigrid; bellemare2013arcade). Works such as WebArena (zhou2023webarena), Mind2Web (deng2023mind2web), and VisualWebArena (koh2024visualwebarena) extend agent capabilities to more realistic, web-based environments. ARE (andrews2025arescalingagentenvironments) further builds upon this line by introducing a mobile-style setting that integrates multiple apps and tools, incorporating time control and a structured verifier to support scalable evaluation. In 3D scenarios, environments emphasize more about embodied interaction, realistic physics and visual perception. More general simulators such as Habitat (savva2019habitat) and ThreeDWorld (gan2020threedworld) support more actions such as navigation and manipulation, while domain-specific worlds develop the specific aspect. For example, MineRL (guss2019minerl) and MineDojo (fan2022minedojo) exploit the flexibility of Minecraft. Besides, Household benchmarks such as ALFRED (shridhar2020alfred) and EmbodiedQA (das2018embodied) further evaluate the grounding function of language in photo-realistic 3D spaces, requiring agents to follow instructions, answer questions, or perform multi-step tasks.
Text-based Environments
Text-based environments emphasize more on reasoning and decision-making from natural language descriptions. TextWorld (cote2018textworld) proposes a framework to generate the text games with different goal-driven challenges. Reddit-RL-simulator (he-etal-2016-deep-reinforcement; chan-king-2018-thread) implements an RL environment for iteratively tracking and recommending popular discussion threads on Reddit. ALFWorld (shridhar2020alfworld) adapts ALFRED tasks into textual form, and JerichoAgentBench (liu2023agentbench) extends the Jericho suite (hausknecht2020interactive) with annotated benchmarks and concrete objectives. Such text environments stress the agent’s ability to infer latent dynamics under partial observability, making them a natural testbed for language-based decision-making.
7 Future Directions
Co-Evolution via Embedded External Tools
As the complexity of task increases, feedback mechanisms must scale in order to avoid sparse, noisy, or misaligned learning signals. Future environments can achieve this by including external tools or modules to serve as verifiers, simulators, compilers, or executable systems into the learning loop. These tools would evolve together with the task generation, providing structured, verifiable feedback and enabling agents to interact with increasingly sophisticated challenges. By integrating embedded tools with automated evaluators such as LLM-based formative signals, environments can better support multi-step, open-ended, and creative problem solving, while also helping to mitigate the Generator-Verifier Asymmetry.
Scaling Through Generator-Verifier Synergy
Future environments can encourage stronger generators to have the ability to decompose complex tasks into smaller subproblems with intermediate solutions, making them tractable for weaker verifiers. This enables scalable supervision in domains where holistic verification is difficult such as creative cultural production and policy-making. In contrast, weaker generators could provide diverse candidate solutions that could be filtered, ranked, or refined by stronger verifiers. By incorporating these co-evolving dynamics into the environment, the asymmetry between generators and verifiers can serve as a catalyst for continuous self-improvement.
Open-Ended, Multi-Agent Environments
Future environments can scale to support large-scale multi-agent interactions, emergent social dynamics, and economic or organizational level simulations, providing rich contexts for studying both collaborative and non-cooperative behaviors in complex environments. In particular, scaling to massively multi-cultural and multi-lingual settings requires scaling environment construction for agents to navigate the subtle semantics of concepts and values that vary across different societies, not only in the textual but also in the multi-modal domain (koh-etal-2024-visualwebarena; huangcultureclip). Such open-ended and interactive scenarios environments foster generalization and strategic planning abilities of agents, equipping LLM agents to better handle complex, real-world challenges beyond isolated task execution.
8 Conclusion
The era of experience makes environments central to the development of LLM agents, positioning them as active producers of experiential data and underscoring the growing need for scaling environments to create a more complex, realistic, and richly interactive world. From a pioneering environment-centric perspective, our survey proposes a unified taxonomy that organizes representative work across the GEF loop (task generation, task execution, and feedback), together with evaluation, implementation, and applications. Besides, we surface key challenges for advancing agent intelligence, including the asymmetry between generators and verifiers and the construction of more open, large-scale multi-agent environments, thereby providing insights for future research on agentic systems.
Limitations
As scaling environments remains an emerging research topic, relatively few studies have explicitly adopted this framing. Thus, we take a broad view of environment and organize representative studies along the GEF loop (task generation, task execution, and feedback) from a pioneering environment-centric perspective. While this broader lens brings in some adjacent lines of work that may not have been explicitly designed from the environment side, our taxonomy is both comprehensive and insightful. Given the rapid pace of agentic research, some of the most recent papers may fall outside this snapshot. We will continue to update this survey as the literature evolves.
Acknowledgments
We thank the anonymous reviewers for their valuable suggestions and comments. Yuchen Huang is supported by the Hong Kong Ph.D. Fellowship Scheme (HKPFS).
Appendix A Conceptual Framework
Following the formalization by gao2025survey, we model the environment $E$ as a partially observable Markov decision process (POMDP). At the beginning of each episode, the environment generates a task $T=(E,I)$, where $I\in\mathcal{I}$ represents user intention drawn from intention space $\mathcal{I}$. The agent $\pi$ interacts with the environment over horizon $T\in\mathbb{N}$, producing an interleaved observation–action trajectory $\tau=(o_{0},a_{0},o_{1},a_{1},\ldots,o_{T})\in(\mathcal{O}\times\mathcal{A})^{T+1}$, where $\mathcal{O}$ and $\mathcal{A}$ denote the observation and action spaces respectively. The environment then evaluates performance and provides feedback $r\in\mathbb{R}^{k}$, which may take the form of step-level signals $r_{\text{step}}^{(t)}=R_{\text{step}}(s_{t},a_{t},T)\in\mathbb{R}$ for $t\in\{0,\ldots,T\}$, where $s_{t}\in\mathcal{S}$ denotes the state of the environment at step $t$, trajectory-level signals $r_{\text{traj}}=R_{\text{traj}}(\tau,T)\in\mathbb{R}^{m}$, or a combination of both $r=f(r_{\text{step}}^{(0:T)},r_{\text{traj}})$ where $f:\mathbb{R}^{T+1}\times\mathbb{R}^{m}\to\mathbb{R}^{k}$. These signals need not be limited to sparse scalar rewards but can encode structured or adaptive assessments reflecting correctness, efficiency, reasoning depth, or long-term outcomes. This Generation-Execution-Feedback (GEF) Loop $\mathcal{L}=(T_{\text{gen}},\text{Exec},\text{Eval}):\mathcal{I}\to\mathcal{T}\times(\mathcal{O}\times\mathcal{A})^{*}\times\mathbb{R}^{k}$, which encompasses task generation, task execution, and feedback, defines the essential mechanics of environments. Repeated iterations drive the accumulation of experience and the progressive evolution of the agent $\pi$.
Appendix B Evaluation Benchmarks
Previous evaluation studies have typically focused on the intelligence of the agents themselves, but there is a lack of direct measurement indicators for aspects such as adaptability to the environment, interactivity, realism, and robustness. Therefore, most environmental assessments are conducted indirectly, usually by observing the performance of intelligent agents to reflect the quality of the environment. For instance, studies such as TaskCraft (shi2025taskcraft) and AgentScaler (fang2025generalagentic) train the agents through the trajectories generated by the interaction between the environment and the agents, thereby evaluating the environment. The stronger performance of the agents is regarded as an indirect indication of higher environmental quality. Initially, direct measurements of the environment are mainly limited to symbolic or textual environments. Bytesized32 (wang2023bytesized32) proposes specific-task text games and evaluates them using automated metrics in terms of fidelity, validity, specification adherence, and winnability. Text2World (hu2025text2world) benchmarks the generation of symbolic world models, using structural similarity for overall evaluation, and capturing more granular features such as action dynamics through component-level F1 scores. Recent studies have begun to extend the direct assessment to more modalities. VidOSC (xue2024learning) explores the dynamic characteristics of open-world environments. WorldScore (duan2025worldscore) proposes a unified framework for evaluating world generation. While WorldPrediction (chen2025worldprediction) focuses on advanced visual reasoning, emphasizing long-term procedural planning and semantic-time abstraction capabilities. Despite these advancements, comprehensive and universal assessment protocols are still scarce, highlighting the need for more generalized and domain-independent methodologies to rigorously and directly evaluate environmental quality beyond the performance metrics of intelligent agents.
Appendix C Key Applications
Recent progress in agentic systems, typically built on state-of-the-art LLM families such as GPT (openai2025introducing), Claude (anthropic2025claude4), Gemini (team2024gemini), LLaMA (touvron2023llama), and Qwen (bai2023qwen), is increasingly driven by interactions with dynamic and multifaceted environments.
Tool-use Environments
Tool-use environments expose APIs and function calls as structured action spaces, and many LLMs now natively support function invocation, thereby extending reasoning with external tools (OpenAIFunctionCalling2023; anthropic2024mcp; wu2025webwalkerbenchmarkingllmsweb; mastouri2025makingrestapisagentready; luo2025mcp; wang2025mcp; fan2025mcptoolbench++).
Coding Environments
Coding environments leverage repositories, test frameworks, and IDE integration to support long-horizon programming. Within these settings, systems such as Qwen3-Coder (qwen3coderblog2024) and Claude 4 (anthropic2025claude4) demonstrate reliability in code editing and debugging. ReTool (feng2025retoolreinforcementlearningstrategic) further integrates code-interpreter execution into the reasoning loop, enabling agents to exhibit code self-correction and adaptive tool selection.
GUI Environments
Web navigation (browser control) environments build on HTML/DOM structures to support tasks such as browsing, form filling, and transactions (lai2024autowebglm; AgenticSeek2024; AutoGPT2023). GUI environments extend these to graphical user interfaces on desktops and mobile devices (qin2025uitars; wang2025opencua; ye2025mobileagentv3; lu2025guiodyssey; liu2024autoglm; sun2022metagui; huang2025guikvefficientguiagents). Beyond these, more comprehensive platforms (wang2025uitars2; xie2024osworld; OpenAICUA2025) integrate terminals, operating systems, applications, and APIs to create more interactive, open-ended, and realistic environments, fostering the development of more advanced computer-use agents (CUAs).
Deep Research Environments
To nurture the next generation of more powerful research agents, deep research environments demand stronger long-context reasoning and more robust retrieval capabilities. Systems such as Gemini 1.5 Pro (google2024geminideepresearch) and OpenAI’s deep research agents (openai2025deepresearch) demonstrate that extended context windows enable sustained, in-depth analysis, while effective retrieval pipelines help mitigate distraction and context dilution.